# Bayes Adaptive Monte Carlo Tree Search for Offline Model-based Reinforcement Learning

### v1

@@ 4.1 The Key Role of Deep Ensembles
The ensemble can also be used for uncertainty quantification. As aforementioned, our algorithm relies on thorough search on the learned world models. Without any constraints on the search process, the learned policy may overfit to an inaccurate model (by overestimating the expected return) and fail in the true MDP. Although the agent could adapt its belief and follow more reliable ensemble members in the Bayesian RL framework, there could be regions in the state-action space where none of the members generalize well, as they are all learned from a static offline dataset. A typical solution is to construct a P-MDP (see Section 3), which lower-bounds the true MDP and discourages the policy from regions where there is large discrepancy between the true and learned world models. We construct the P-MDP by modifying each reward estimation $r$ into $\tilde{r}$: ($\mu_{\theta}(s,a)=\sum_{i=1}^{K}b(\theta)(i)\mu_{\theta}^{i}(s,a)$)
$\tilde{r}(s,a,r,b(\theta))=r-\lambda\sqrt{\sum_{i=1}^{K}b(\theta)(i)(\sigma_{\theta}^{i}(s,a)^{2}+\mu_{\theta}^{i}(s,a)^{2})-\mu_{\theta}(s,a)^{2}}$ (5)
The reward penalty is weighted by a hyperparameter $\lambda>0$ and corresponds to the standard deviation (std) of the mixture of Gaussian dynamics models, where $\mu_{\theta}^{i}$ and $\sigma_{\theta}^{i}$ are the mean and std from the ensemble member $i$. This penalty design combines epistemic and aleatoric model uncertainty and has been shown to be successful at capturing errors in predicted dynamics (Lu et al. (2022)) [^3] In the original literature (Lakshminarayanan et al. (2017); Lu et al. (2022)), the ensemble is treated as a uniformly-weighted mixture model, i.e., $b(\theta)(i)=1/K,\ (i=1,\cdots K)$, since their belief is not adaptive. Equation (5) fits into the Bayesian RL framework by adapting the belief $b(\theta)$, which is part of our novelty..
@@
4.2 Bayes Adaptive MCTS in Continuous State and Action Spaces
BAMCP (Guez et al. (2013)) has been successful in solving large-scale BAMDPs, as detailed in Appendix A, but it is limited to scenarios with discrete state and action spaces. In this subsection, we introduce an online planning method to approximate the Bayes-optimal policy at a decision point $(s,h)$. ($h$ denotes the transition history that ends at $s$.) In particular, this method can be used to solve BAMDPs with continuous states and actions. Specifically, we adopt double progressive widening (DPW, Couëtoux et al. (2011); Auger et al. (2013)), which maintains a finite list of chance nodes to be searched at each decision point and incrementally adds a new chance node to the list based on the visitation counts. For a node $(s,h)$, a new action $a$ is sampled (with the current policy) and added to its children set $C((s,h))$, if $\left\lfloor{N((s,h))^{\alpha}}\right\rfloor\geq|C((s,h))|$, where $\alpha\in(0,1)$ is a hyperparameter that controls the growth rate and $N$ denotes the visitation counts. Otherwise, an action will be sampled from $C((s,h))$ according to the UCT (Kocsis & Szepesvári (2006)) rule. Similarly, to handle the infinitely many possible transitions, a new next state $s^{\prime}$ is added to the children set $C((s,h),a)$ only if $\left\lfloor{N((s,h),a)^{\beta}}\right\rfloor\geq|C((s,h),a)|$ ($\beta\in(0,1)$). Otherwise, the least visited child in $C((s,h),a)$ will be selected as the next state. With DPW, the sets of possible actions or next states for search are all finite, allowing deep search as in discrete scenarios, and the more promising states and actions (with higher $N$) have more subsequent branches, thereby reducing corresponding estimation uncertainty. However, directly integrating DPW and BAMCP (i.e., Algorithm 3) likely cannot solve BAMDPs with continuous state and action spaces. As introduced in Appendix A, BAMCP relies on root sampling, which samples dynamics functions (from $b(\theta)$) only at the root node and follows a specific dynamics function throughout a simulation procedure. However, the rationale of root sampling (i.e., Lemma A.1) does not hold when applying DPW [^4] The last equality of Eq. (6) does not hold, since $\tilde{b}(\theta|has^{\prime})\propto\tilde{b}(\theta|ha)\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)\neq\tilde{b}(\theta|ha)\mathcal{P}_{\theta}(s^{\prime}|s,a)$. $\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)$ represents the state transition distribution when applying DPW, which differs from the true distribution $\mathcal{P}_{\theta}(s^{\prime}|s,a)$, as dictated by the DPW rule.. Also, BAMCP assumes that the reward function is certain rather than a function of $\theta$.
@@ 4.2 Bayes Adaptive MCTS in Continuous State and Action Spaces
On the other hand, Polynomial Upper Confidence Trees (PUCT, Auger et al. (2013)), based on DPW, is a provably consistent planning method for solving MDPs with infinite-scale state/action spaces and arbitrarily stochastic transition kernels. In this case, we propose casting BAMDPs into MDPs (i.e., $\mathcal{M}^{+}$ defined in Section 2) and solving them with PUCT. The pseudo code is shown as Algorithm 1. Ideally, as the number of simulations $E\rightarrow\infty$, PUCT can find a near-optimal solution of $\mathcal{M}^{+}$, which is also a near Bayes-optimal solution for the true environment. Each simulation follows a path that begins at the root node and ends at a leaf node, utilizing progressive widening for sampling both actions and next states, as illustrated in the ActionPW and StatePW procedures. Compared to PUCT, the significant modifications include: (1) replacing $\langle\mathcal{S},\mathcal{P},\mathcal{R}\rangle$ with their extended definitions in BAMDPs, $\langle\mathcal{S}^{+},\mathcal{P}^{+},\mathcal{R}^{+}\rangle$, and (2) applying reward penalties to account for model uncertainty. To be specific, in StatePW, $r$ and $s^{\prime}$ are sampled from a mixture of ensemble members, which is a practical implementation of sampling from $\mathcal{R}^{+}$ and $\mathcal{P}^{+}$ as shown in Eq. (1). After receiving the transition $(s,a,r,s^{\prime})$, the belief vector $b(\theta)$ is updated to $b^{\prime}(\theta)$ following Eq. (4), finishing the transition in $\mathcal{S}^{+}$ from $(s,b(\theta))$ to $(s,b^{\prime}(\theta))$. Moreover, the reward $r$ is adjusted with a penalty term as defined in Eq. (5), which is then used to calculate the return $R$. As aforementioned, applying such a reward penalty can effectively mitigate the issue of model exploitation.
@@
4.3 The Overall Framework: Search-based Policy Iteration
In this subsection, we present a framework (inspired by MuZero [^6] The novelty of our algorithm in comparison to MuZero is detailed in Section 3.) that integrates continuous BAMCP (i.e., Algorithm 1) into policy improvement and policy evaluation. By iteratively running these procedures, we can approach a near Bayes-optimal policy, i.e., $\pi$, that can be directly referred to during execution in the true environment. The pseudo code of the overall framework is shown as Algorithm 2. For efficiency, a learner and a number of actors execute in parallel, reading from and sending data to the replay buffer $\mathcal{D}$ respectively. The actors update their copies of policy and value functions every $E_{l}$ learner steps.
In particular, each actor would interact with the learned world models to sample trajectories $\tau$ [^7] In Algorithm 2, $\hat{\rho}(s)$ denotes the empirical distribution of initial states in the offline dataset $\mathcal{D}_{\mu}$. At each time step of the trajectory, a Search procedure (defined in Algorithm 1) is executed using the actor’s current copy of $\pi$ and $V$ for planning at the current decision point. The search result $\pi_{\text{ret}}$ is then used to indicate the action choice, i.e., $a\sim\pi_{\text{ret}}(\cdot|(s,h))$. The subsequent transition process follows a BAMDP, where $r\sim\mathcal{R}^{+}(\cdot|(s,h),a)$, $s^{\prime}\sim\mathcal{P}^{+}(\cdot|(s,h),a)$, and the belief $b({\theta})$ adapts with experience, as described in StatePW. The collected trajectories are used in the learning process, where $\pi$ is trained to mimic the search result $\pi_{\text{ret}}$ by minimizing a cross-entropy loss (i.e., $\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})$), while $V$ is trained to predict the $n$-step pessimistic return $z$ [^8] In the definition of $z$, $\tilde{r}^{(0)}=\tilde{r}$ is the pessimistic reward associated with $(s,h)$ and the subscript $(i)$ denotes that the corresponding quantity is collected in $i$ time steps within that trajectory. As noted in (Hubert et al. (2021)), $\pi_{\text{ret}}$ improves $\pi$ at each decision point, so repeatedly applying continuous BAMCP to obtain $\pi_{\text{ret}}$ and projecting the search results to the parameter space of $\pi$ (by minimizing $\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})$) represents the policy improvement process. Meanwhile, value backups to update the Q estimates (i.e., $R\leftarrow V((s^{\prime},hars^{\prime}))$ in Algorithm 1) and the temporal difference learning for the value function $V$ constitute the policy evaluation process. In this way, the search algorithm is used as a powerful improvement operator to iteratively enhance the performance of the learned policy $\pi$.
Algorithm 2 Search-based Policy Iteration
@@ 4.3 The Overall Framework: Search-based Policy Iteration
Initialize $\pi$ and $V$, $\mathcal{D}\leftarrow\emptyset$
@@
         $\{\tau_{i}\}_{i=1}^{B}\sim\mathcal{D}$
         $\min_{\pi}\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})$, $\min_{V}\mathcal{L}(V,\{\tau_{i}\}_{i=1}^{B})$
[…]
         Update $\pi,V$ in Actor if $e\%E_{l}==0$
@@
$\mathcal{L}(\pi,\{\tau_{i}\}_{i=1}^{B})=-\sum_{((s,h),\pi_{\text{ret}})}\pi_{\text{ret}}^{T}\log\pi(\cdot|(s,h))/(BT)$
$\mathcal{L}(V,\{\tau_{i}\}_{i=1}^{B})=\sum_{((s,h),z)}(V(s,h)-z)^{2}/(BT)$,
$z=\tilde{r}^{(0)}+\gamma\tilde{r}^{(1)}+\cdots+\gamma^{n-1}\tilde{r}^{(n-1)}+\gamma^{n}v_{\text{ret}}^{(n)}$
@@
         $s\leftarrow\hat{\rho}_{0}(\cdot)$, $h\leftarrow s$, $\tau\leftarrow[]$
         $b(\theta)\leftarrow[1/K,\cdots,1/K]$
         for $t=1\cdots T$ do
@@ 5 Evaluation
To evaluate the effectiveness of each component in our algorithm design, we introduce three variants: (1) BA-MBRL leverages learned world models as surrogate simulators, applying reward penalties to collected transitions and using standard online RL algorithms (e.g., SAC (Haarnoja et al. (2018))) to learn a policy. While following existing offline MBRL methods, it models the problem as a BAMDP (rather than an MDP), with environment transitions defined by Equations (1) and (4) and the reward penalty by Equation (5), and is designed to evaluate the effectiveness of Bayesian RL. (2) BA-MCTS builds on BA-MBRL by introducing Continuous BAMCP (Algorithm 1) to plan at decision points, rather than inferring directly from the policy, to generate trajectories for downstream SAC, demonstrating the impact of deep search on policy learning. (3) BA-MCTS-SL, described in Algorithm 2, replaces the policy learning algorithm in BA-MCTS from policy gradient methods (as in SAC) with supervised learning (SL), allowing us to compare which approach offers a more efficient policy update mechanism, particularly for continuous control tasks.
Data Type Environment BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) Optimized COMBO MOReL MOPO
random HalfCheetah 29.20 $\pm$ 2.00 36.23 $\pm$ 1.04 32.76 $\pm$ 1.16 31.7 38.8 25.6 35.4
random Hopper 33.83 $\pm$ 0.10 31.56 $\pm$ 0.12 31.47 $\pm$ 0.03 12.1 17.9 53.6 11.7
random Walker2d 21.89 $\pm$ 0.07 21.59 $\pm$ 0.32 21.45 $\pm$ 0.53 21.7 7.0 37.3 13.6
medium HalfCheetah 70.47 $\pm$ 3.52 75.84 $\pm$ 3.81 56.54 $\pm$ 5.20 45.7 54.2 42.1 42.3
medium Hopper 97.75 $\pm$ 7.09 96.70 $\pm$ 14.0 98.25 $\pm$ 3.42 69.3 97.2 95.4 28.0
medium Walker2d 82.24 $\pm$ 1.85 74.73 $\pm$ 3.25 75.41 $\pm$ 4.17 79.7 81.9 77.8 17.8
med-replay HalfCheetah 61.16 $\pm$ 1.60 65.45 $\pm$ 0.81 62.50 $\pm$ 0.18 58.0 55.1 40.2 53.1
med-replay Hopper 106.3 $\pm$ 0.13 101.8 $\pm$ 3.46 93.91 $\pm$ 4.25 90.8 89.5 93.6 67.5
med-replay Walker2d 92.13 $\pm$ 5.13 95.06 $\pm$ 2.11 97.54$\pm$ 1.93 65.8 56.0 49.8 39.0
med-expert HalfCheetah 80.53 $\pm$ 6.63 76.16 $\pm$ 10.3 90.52 $\pm$ 4.13 104.2 90.0 53.3 63.3
med-expert Hopper 112.2 $\pm$ 0.29 108.3 $\pm$ 0.22 107.8 $\pm$ 0.37 105.8 111.1 108.7 23.7
med-expert Walker2d 107.7 $\pm$ 0.82 110.0 $\pm$ 1.74 84.71 $\pm$ 0.87 97.1 103.3 95.6 44.6
Average Score 74.62 74.45 71.06 65.16 66.83 64.42 36.67

### v4

@@ 4.1 The Key Role of Deep Ensembles
Although the agent could adapt its belief and follow more reliable ensemble members in the Bayesian RL framework, there could be regions in the state-action space where none of the members generalize well, as they are all learned from a static offline dataset. A typical solution for such uncertainty is to construct a pessimistic MDP, which discourages the agent from exploiting regions with high uncertainty. Another use of the deep ensemble is to construct a Pessimistic Bayes-Adaptive MDP (P-BAMDP).
In particular, we apply a reward penalty to each transition $(s,a,r)$ and get a penalized reward $\tilde{r}$ as follows:
$r-\lambda\cdot\text{std}\left[r^{i}+\gamma\mathbb{E}_{\begin{subarray}{c}s^{\prime i}\sim\mathcal{P}_{\theta}^{i},a^{\prime}\sim\pi(s^{\prime i})\end{subarray}}[Q_{\psi^{-}}(s^{\prime i},a^{\prime})]\right]_{i=1}^{K}$ (5)
Here, $s^{\prime i}\sim\mathcal{P}_{\theta}^{i}(\cdot|s,a)$ and $r^{i}=\mathbb{E}[\mathcal{R}_{\theta}^{i}(\cdot|s,a)]$ denote the next state and expected reward from the $i$-th ensemble member. $\pi$ and $Q_{\psi^{-}}$ represent the policy and target Q-network. This penalty term is the standard deviation of the one-step lookahead Q-value targets, computed across all ensemble members. The underlying intuition is that a high variation in these predictions indicates a high degree of uncertainty at this region. As demonstrated by Sun et al. (2023), employing this penalty to define a pessimistic MDP yields an agent performance estimate that more closely matches real-world outcomes, compared to penalties based solely on ensemble disagreement in next-state prediction, as proposed in Yu et al. (2020); Kidambi et al. (2020).
4.2 Continuous BAMCP
BAMCP (Guez et al., 2013) has been successful in solving large-scale BAMDPs, as detailed in Appendix A, but it is limited to scenarios with discrete state and action spaces. In this subsection, we introduce a novel planning method to approximate the Bayes-optimal policy at a decision point $(s,h)$ ($h$ denotes the transition history that ends at $s$), which can be used to solve BAMDPs with continuous states/actions and stochastic transition kernels.
Double Progressive Widening (DPW): DPW (Couëtoux et al., 2011; Auger et al., 2013) is a technique to extend the use of MCTS to continuous state and action spaces. Instead of exploring all possible actions and next states, DPW maintains a finite list of options to search at each decision node, incrementally adding new options to the list based on the visitation counts of that decision node. To be specific, for a node $(s,h)$, a new action $a$ is sampled (with the current policy $\pi$) and added to its children set $C((s,h))$, if $\left\lfloor{N((s,h))^{\alpha}}\right\rfloor\geq|C((s,h))|$, where $\alpha\in(0,1)$ is a hyperparameter that controls the growth rate and $N$ denotes the visitation counts of $(s,h)$. Otherwise, an action is selected from existing options in $C((s,h))$ according to the UCT (Kocsis and Szepesvári, 2006) rule. Similarly, to handle the infinitely many possible state transitions in stochastic environments, a new next state $s^{\prime}$ is added to the children set $C((s,h),a)$ only if $\left\lfloor{N((s,h),a)^{\beta}}\right\rfloor\geq|C((s,h),a)|$ ($\beta\in(0,1)$). Otherwise, the least visited child in $C((s,h),a)$ will be selected as the next state. With DPW, the sets of possible actions or next states to explore are finite, allowing deep tree search as in discrete scenarios.
Integration of DPW and BAMCP: Directly combining DPW and BAMCP (i.e., Algorithm 3) cannot solve BAMDPs with continuous state and action spaces. As introduced in Appendix A, BAMCP relies on root sampling, which samples dynamics functions only at the root node and follows a specific dynamics function throughout a simulation rollout. However, the rationale of root sampling (i.e., Lemma A.1) does not hold when applying DPW [^1] The last equality of Eq. (6) does not hold, since $\tilde{b}(\theta|has^{\prime})\propto\tilde{b}(\theta|ha)\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)\neq\tilde{b}(\theta|ha)\mathcal{P}_{\theta}(s^{\prime}|s,a)$. $\widetilde{\mathcal{P}}_{\theta}(s^{\prime}|s,a)$ represents the distribution of next states when applying DPW, which differs from the true distribution $\mathcal{P}_{\theta}(s^{\prime}|s,a)$, as dictated by the DPW rule.. As an alternative design, Polynomial Upper Confidence Tree (PUCT) (Auger et al., 2013), built upon DPW, is a provably consistent planning method for solving MDPs with infinite-scale state and action spaces and highly stochastic transition dynamics. Thus, we propose to recast P-BAMDPs as MDPs and solve them using PUCT, as a novel Bayes planning method for continuous state and action spaces.
@@ 4.2 Continuous BAMCP
Proposed algorithm – Continuous BAMCP: In Algorithm 1, each simulation follows a path from the root node to an unvisited node, utilizing progressive widening when sampling actions or next states, as detailed in the ActionPW and StatePW procedures. Compared to PUCT, the main modifications include: (1) replacing $\langle\mathcal{S},\mathcal{P},\mathcal{R}\rangle$ in MDPs with their extended definitions in BAMDPs, i.e., $\langle\mathcal{S}^{+},\mathcal{P}^{+},\mathcal{R}^{+}\rangle$, and (2) applying reward penalties to account for model uncertainty. To be specific, in StatePW, $r$ and $s^{\prime}$ are sampled from the distribution predicted by all ensemble members, which is a practical implementation of sampling from $\mathcal{R}^{+}$ and $\mathcal{P}^{+}$ as outlined in Eq. (1). After receiving the transition $(s,a,r,s^{\prime})$, the belief vector $b(\theta)$ is updated to $b^{\prime}(\theta)$ following Eq. (4), finishing the transition in $\mathcal{S}^{+}$ from $(s,b(\theta))$ to $(s^{\prime},b^{\prime}(\theta))$. The transition history $h$ is then updated to $h^{\prime}=hars^{\prime}$. Secondly, in Simulate, a penalized reward $\tilde{r}$ (i.e., Eq. (5)) is used, which effectively mitigates overfitting to inaccurate ensemble members by incorporating pessimism, as discussed in Section 4.1.
The following theorem establishes that our planner is a reliable and consistent solver for P-BAMDPs defined in Section 4.1. The detailed proof is provided in Appendix B. This consistency guarantee is crucial, as it ensures that with sufficient computation, our planner’s output converges to the true optimal value within the pessimistic model, providing a reliable value target for our overall learning framework.
Theorem 4.1.
Let $\hat{V}(z_{0})$ be the value of the root node $z_{0}$ estimated by Continuous BAMCP after $E$ simulations. Under the regularity and value function conditions defined in Appendix B.1, the error of this estimate with respect to the P-BAMDP’s optimal value $V^{*}_{p}(z_{0})$ is bounded as follows, exponentially surely in $E$:
$\left|\hat{V}(z_{0})-V^{*}_{p}(z_{0})\right|\leq C_{0}E^{-\gamma_{0}}+\gamma^{d_{max}}\epsilon_{V},$
where $\gamma_{0},C_{0}>0$ are constants, $\gamma$ is the discount factor, $d_{max}$ is the maximum search depth, and $\epsilon_{V}$ is the uniform bound on the value approximation error at the leaf nodes.
@@
Algorithm 2 BA-MCTS
@@
Initialize the actor $\pi$ and critic $Q_{\psi}$, $\mathcal{D}\leftarrow\emptyset$
@@
   $\{(s,h,\pi_{\text{ret}},\tilde{r},s^{\prime},h^{\prime})_{i}\}_{i=1}^{B}\sim\mathcal{D}$
   Update $\pi$ and $Q_{\psi}$ using SAC
   $e\mathrel{{+}{=}}1$
   Update $\pi$ and $V$ (defined on $Q_{\psi}$) in Actor if
   $e\%E_{l}==0$
@@
   Sample $s$ from $\mathcal{D}_{\mu}$, $h\leftarrow s$, $\tau\leftarrow[]$
   Obtain the prior $b(\theta)$ at $s$
   for $t=1\cdots H$ do
@@
4.3 Overall Framework: BA-MCTS
Now, we present how to integrate planning outcomes of continuous BAMCP into a policy iteration framework, to obtain a near Bayes-optimal policy, i.e., $\pi$, that can be directly referred to during execution. The pseudo code is shown as Algorithm 2. For efficiency, a learner and a number of actors execute in parallel, reading from and sending data to the replay buffer $\mathcal{D}$ respectively. The actors update their copies of policy and value functions every $E_{l}$ learner steps.
Each actor interacts with the learned world models to sample trajectory segments $\tau$. The starting states of these segments are sampled from the provided dataset $\mathcal{D}_{\mu}$. Notably: (1) the segment length $H$ is kept short to minimize error accumulation when interacting with the learned world models; and (2) the prior $b(\theta)$ at a starting state $s$ is obtained by performing the Bayesian posterior update (i.e., Eq. (4)) on the offline trajectory in $\mathcal{D}_{\mu}$ containing $s$. These belief updates are reliable, as the offline trajectories are from the real environment. At each decision node $(s,h)$ of the segment, a Search procedure (defined in Algorithm 1) is executed. The search result $\pi_{\text{ret}}$ then indicates the action choice, i.e., $a\sim\pi_{\text{ret}}(\cdot|(s,h))$. Given $a$, the transition to the next state follows a BAMDP, where $r\sim\mathcal{R}^{+}(\cdot|(s,h),a)$, $s^{\prime}\sim\mathcal{P}^{+}(\cdot|(s,h),a)$, and the belief $b({\theta})$ is adapted with the new transition.
With the collected segments $\tau$, we update the policy $\pi$ and value function $V$. Our primary algorithm, BA-MCTS, employs an off-policy RL method: SAC (Haarnoja et al., 2018). The value function (critic) is updated using standard temporal difference learning based on the sampled transitions, while the policy (actor) is updated using the policy gradient derived from the critic. Additionally, we investigate an alternative approach, BA-MCTS-SL, where the policy is updated via supervised learning to mimic the search result $\pi_{\text{ret}}$ by minimizing a cross-entropy loss: $\mathcal{L}(\pi,\{(s,h,\pi_{\text{ret}})_{i}\}_{i=1}^{B})=-\sum_{((s,h),\pi_{\text{ret}})}\pi_{\text{ret}}^{T}\log\pi(\cdot|(s,h))/B$. We compare these two approaches in our evaluation. In both methods, we seek to distill the superior, non-parametric search outcomes from Continuous BAMCP into the parameterized policy $\pi$.
@@ 5 Evaluation
Section 5.1 discusses the benefit of Bayesian model adaptation. Section 5.2 presents comprehensive benchmarking results on D4RL MuJoCo tasks. Section 5.3 details our ablation studies. Finally, Section 5.4 validates our algorithm’s applicability on a challenging Tokamak control task.
Table 1: Comparison of the prediction errors for next states and rewards in imaginary rollouts, with and without Bayesian belief adaptation. Ground truth for the imaginary rollouts is obtained by replaying the action sequences in the real simulators.
Data Type Environment Prediction Error on Next State Prediction Error on Reward Overall Prediction Error
Adaptive Uniform Adaptive Uniform Adaptive Uniform
random HalfCheetah 0.339 (.021) 0.342 (.017) 0.016 (.001) 0.016 (.001) 0.177 (.011) 0.179 (.009)
random Hopper 0.232 (.016) 0.189 (.008) 0.012 (.001) 0.022 (.008) 0.122 (.008) 0.106 (.008)
random Walker2d 62.10 (11.0) 112.0 (19.0) 1.364 (.215) 4.908 (.708) 31.73 (5.61) 58.44 (9.83)
medium HalfCheetah 1.387 (.040) 1.355 (.038) 0.057 (.001) 0.061 (.001) 0.722 (.021) 0.708 (.019)
medium Hopper 0.429 (.033) 0.570 (.035) 19.03 (8.58) 68.24 (11.4) 9.727 (4.30) 34.40 (5.69)
medium Walker2d 34.01 (.556) 34.26 (.142) 24.38 (4.31) 113.1 (3.47) 29.19 (1.89) 73.69 (1.78)
med-replay HalfCheetah 0.677 (.027) 0.707 (.036) 0.115 (.005) 0.114 (.006) 0.396 (.016) 0.410 (.021)
med-replay Hopper 0.170 (.022) 0.212 (.054) 1.177 (.420) 3.320 (1.14) 0.674 (.221) 1.766 (.558)
med-replay Walker2d 66.83 (7.39) 44.61 (1.34) 27.21(3.24) 60.52 (3.17) 47.02 (5.17) 52.57 (2.25)
med-expert HalfCheetah 1.423 (.054) 1.467 (.046) 0.071 (.002) 0.074 (.002) 0.747 (.028) 0.771 (.024)
med-expert Hopper 3.665 (2.56) 18.27 (2.09) 71.46 (79.9) 384.7 (37.6) 37.56 (41.2) 201.5 (19.4)
med-expert Walker2d 55.69 (3.71) 51.89 (.481) 121.9 (12.6) 186.3 (3.86) 88.77 (8.18) 119.1 (2.08)
5.1 The Benefit of Belief Adaptation
The core idea of belief adaptation (using the BAMDP framework detailed in Section 4.1) is that by dynamically adjusting the belief over each ensemble member based on observed transitions, the agent can rely more on models that are accurate for the current trajectory, leading to better planning. Our empirical results show that belief adaptation not only improves model prediction accuracy, as shown in Table 1, but also translates to better final policy performance on offline RL tasks, as shown in Tables 2 and 4. Specifically, Table 1 highlights that the adaptive ensemble achieves more accurate predictions for next states and rewards in 10 out of 12 benchmarking tasks, compared to treating each ensemble member uniformly as in previous offline MBRL methods. This suggests that Bayesian belief adaptation leads to a more accurate world model for planning, which is a key factor underlying observed performance improvements. A more detailed study of this aspect, including a visualization of the belief adaptation process within rollouts, is provided in Appendix F.1.
5.2 Benchmarking Results on D4RL MuJoCo
We evaluate our algorithm on a widely-used offline RL benchmark: D4RL MuJoCo (Fu et al., 2020), comparing its performance with state-of-the-art (SOTA) model-based/model-free methods, and search-based approaches.
Model-Based Methods. In Table 2, BA-MCTS achieves the highest average score across all twelve D4RL MuJoCo tasks when compared against a range of SOTA model-based baselines. This result confirms that the combination of a principled Bayesian treatment of uncertainty and a powerful search-based planner leads to superior performance.